OpenStudy (anonymous):

5 times the square root of 700?

OpenStudy (anonymous):

5 x SQR of 700 = 132.288

OpenStudy (isaiah.feynman):

\[5\sqrt{700} = 5\sqrt{7\times100}=50\sqrt{7}\]

OpenStudy (anonymous):

^ I'd go with that answer mine is simplified

OpenStudy (anonymous):

thank you guys